A jeweler wants to melt a 1.25 kg ingot of silver metal by heating it from 20.0oC to its melting point of 961oC, before pouring it into moulds. Calculate the energy required to carry out the procedure if the specific heat of solid silver is 233 J kg-1 K-1 and the latent heat of fusion of silver is 1.11 x 103 J kg-1.
1
Expert's answer
2019-10-24T09:35:43-0400
"Q=mc(T_2-T_1)+mL"

"Q=1.25(233(961-20)+1110)=275000\ J=275\ kJ"